8.Given the demand function Pd = 35 − Q2

and the supply function Ps = 3 + Q2.


Calculate the producer surplus for the supply function.

 

1.      0,0

2.      33,3

3.      69,0

4.      42,7


1
Expert's answer
2021-07-06T14:43:04-0400

Question

Given the demand function Pd =35-Q2 and the supply function Ps =3+Q2 .

Calculate the producer surplus for the supply function.

1.0,0

2.33,3

3.69,0

4.42,7

Solution;

At equilibrium, Pd =Ps

35-Q2 =3+Q2

35-3 =2Q2

16=Q2

Q=4

Using Q=4,Pd =Ps=19

Producer surplus is given by;

P*Q -"\int"0Q(PsdQ)

19(4) -"\int"04(3+Q2)dQ

76-[(3Q+Q3/3)|04

76-[3(4) +43/3]

76-100/3

42.666666...

42.7

Answer;

4. 42,7